# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ 4 Estimating Latent Subspaces from Offline Data
To minimize the regret, one must learn $\bm{\beta}_{\star}$ online. However, it is much easier to search among all $\bm{\theta}\in\mathbb{R}^{d_{K}}$ than to search among all possible $\bm{\beta}\in\mathbb{R}^{d_{A}}$ since typically, $d_{K}\ll d_{A}$. So, it will help to learn some $\hat{\textbf{U}}^{\top}\approx\textbf{U}_{*}^{\top}:\mathbb{R}^{d_{A}}\to\mathbb{R}^{d_{K}}$ offline so that for any estimate $\hat{\bm{\beta}}_{t}$ of $\bm{\beta}_{\star}$, $\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{t}$ is an estimate of $\hat{\bm{\theta}}_{t}\in\mathbb{R}^{d_{K}}$. This amounts to learning a subspace of the feature space from logged bandit data. We therefore provide a method for Subspace estimation from Offline Latent bandit Data (SOLD), presented as Algorithm 1.
@@
3:  Estimate reward parameters $\hat{\bm{\beta}}_{n,i}\leftarrow\textbf{V}_{n,i}^{-1}\textbf{b}_{n,i}$, where $\textbf{V}_{n,i}=\mu\textbf{I}+\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)\phi(x,a)^{\top}$ and $\textbf{b}_{n,i}=\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)r$ for $i=1,2$.
4:  Compute $\textbf{M}_{n}=\frac{1}{2}(\hat{\bm{\beta}}_{n,1}\hat{\bm{\beta}}_{n,2}^{\top}+\hat{\bm{\beta}}_{n,2}\hat{\bm{\beta}}_{n,1}^{\top})$ and compute $\overline{\textbf{M}}_{N}:=\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}$.
5:  Compute matrices $\overline{\textbf{D}}_{N,i}:=\frac{1}{N}\sum_{n=1}^{N}(I-\mu\textbf{V}_{n,i}^{-1})$ for $i=1,2$.
6:  Obtain the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$, denoting them by $\hat{\textbf{U}}$.
@@
We split each trajectory to obtain two independent estimators $\hat{\bm{\beta}}_{n,1},\hat{\bm{\beta}}_{n,2}$ of $\bm{\beta}$, compute the outer products $\hat{\bm{\beta}}_{n,1}^{\top}\hat{\bm{\beta}}_{n,2}$ for each trajectory, and try to obtain the top $d_{K}$ eigenvectors of the mean outer product across trajectories. We want the mean outer product to concentrate to $\mathbb{E}[\bm{\beta}\bm{\beta}^{\top}]=\textbf{U}_{\star}\mathbb{E}[\bm{\theta}\bm{\theta}^{\top}]\textbf{U}_{\star}^{\top}$, which has the same $d_{K}$-dimensional span as $\textbf{U}_{\star}$.
However, there is a wrinkle here. We cannot simply take the top $d_{K}$ eigenvectors of the mean outer product $\overline{\textbf{M}}_{N}$. Note that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}\textbf{U}_{\star}\bm{\theta}_{n}\bm{\theta}_{n}^{\top}\textbf{U}_{\star}^{\top}\textbf{D}_{n,2}]$. To separate $\mathbb{E}[\bm{\beta}_{n}\bm{\beta}_{n}^{\top}]$ from this, we need $\textbf{D}_{n,1},\bm{\beta}_{n},\textbf{D}_{n,2}$ to be independent. If $\pi_{b}$ does not use $\bm{\theta}$ and contexts are generated independently of each other and of $\bm{\theta}$, then this is satisfied. In fact, we show in the lemma below that if any of these three conditions is violated, then it is in fact impossible to determine the latent subspace $\textbf{U}^{\star}$ using any method, even with infinitely many infinitely long trajectories.
@@
there exists a pair of linear latent contextual bandits with orthogonal latent subspaces satisfying the condition, and a behavior policy $\pi_{b}$ so that the offline data distributions are indistinguishable and cover all $(x,a)$ pairs with probability at least $1/4$. Since the latent subspaces are orthogonal, an action that gives the maximum reward on one latent bandit gives reward 00 on the other.
@@
As an example, this assumption is satisfied in practice when the offline data comes from randomized controlled trials or A/B testing, which are common sources of offline datasets. [^3] Even when the assumption is not satisfied in practice, Algorithm 1 can learn a good subspace whenever $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ has eigenspace close to the span of $\textbf{U}_{\star}$, e.g. when high-reward actions contribute heavily to $\textbf{D}_{n,i}$. Now, let the covariance matrix of $\bm{\theta}$ be $\Lambda$ and let its mean be $\mu_{\theta}$. Then we have that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}]\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}\mathbb{E}[\textbf{D}_{n,2}]$. So, we still cannot merely consider the top $d_{K}$ eigenvectors of $\overline{\textbf{M}}_{N}$ without accounting for $\textbf{D}_{n,1}$. Intuitively, $\textbf{D}_{n,1}$ captures the fact that we cannot estimate rewards of actions that we have not taken in a short trajectory. We therefore construct correction matrices $\overline{\textbf{D}}_{N,i}$ and use them to "neutralize" the effect of unseen actions from the short trajectories. In particular, $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is an estimator for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Crucially, this allows us to aggregate information across many trajectories to overcome the challenge of learning from short trajectories. We can now take the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to estimate the subspace determined by $\textbf{U}_{\star}$. However, to give guarantees for our method, we must first make a coverage assumption:
@@
Rewards $|r_{n,h}|\leq R$ for all $n,h$, $\|\phi(x,a)\|_{2}\leq 1$ and $\|\bm{\beta}\|_{2}\leq R$. Also, $\lambda_{A}:=\min_{i=1,2}\lambda_{\min}(\mathbb{E}[\textbf{D}_{n,i}])>0$ and $\lambda_{\theta}:=\frac{1}{R^{2}}\lambda_{\min}(\Lambda)>0$.
Intuitively, $\lambda_{A}$ measures coverage along actions, while $\lambda_{\theta}$ measures coverage along latent states $\bm{\theta}$. Both must be non-zero to expect satisfactory estimation of the subspace. We can then use confidence bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ to give a data-dependent confidence bounds $\Delta_{\operatorname{off}}$ for the projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, as in Theorem 2 below. In one instantiation, Propositions 1 and 2 in Appendix C derive simple data-dependent bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ respectively. Under this choice, we can control the growth of $\Delta_{\operatorname{off}}$ in terms of the unknown problem parameters, shown at the end of Theorem 2.
@@
Let $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ and $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}\leq\Delta_{D}$ for $i=1,2$ with probability $31-\delta/$ 3 each. Then, with probability $1-\delta$, $\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}\leq\Delta_{\operatorname{off}}$, where for $B_{D}=\|\overline{\textbf{D}}_{N}^{-1}\|_{2}$ and $\hat{\lambda}:=\lambda_{d_{K}}(\overline{\textbf{M}}_{N})-\lambda_{d_{K}+1}(\overline{\textbf{M}}_{N})$,
$\Delta_{\operatorname{off}}=\textstyle{\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\left(\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}\right)(R^{2}+\Delta_{M})\Delta_{D}+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}\right)}$
@@
Estimating $d_{K}$ from offline data.
Since our estimator $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is approximately rank-$d_{K}$, the number of significant eigenvalues of the estimator is a principled heuristic for determining $d_{K}$.
@@
Naively performing PCA on the raw rewards or on single reward estimates $\hat{\beta}_{n}$ can lead to erroneous subspaces – as while the PCA target is linear-algebraically similar to $\overline{\textbf{M}}_{N}$, it is statistically different. The PCA target (e.g. $\mathbb{E}[\hat{\beta}_{n}\hat{\beta}_{n}^{\top}]$) is typically full rank due to the variance of the per-reward noise $\epsilon$. On the other hand, PMF (Mnih and Salakhutdinov, 2007b) offers neither confidence bounds on the estimated subspace, nor a principled method for determining $d_{K}$.
5 Using Offline Data to Speed Up Online Learning with LOCAL-UCB
Here, we motivate and describe LOCAL-UCB, a natural algorithm that accelerates LinUCB with offline data. The core idea is sharpening optimism by being optimistic over the intersection of two confidence sets – one obtained using offline data and another purely from online data.
To geometrically motivate our update rule, say $t$ online interactions have been completed, yielding a dataset $\mathcal{D}_{t}$ of $(x_{s},a_{s},r_{s})$ tuples. From Theorem 2, SOLD gives an estimate and a confidence set for the projection matrix $\textbf{U}_{\star}\textbf{U}_{\star}^{\top}$. Let’s call the set $\mathcal{C}_{U,\operatorname{off}}$. For each projection $\textbf{U}\textbf{U}^{\top}\in\mathcal{C}_{U,\operatorname{off}}$, we have a $d_{K}$-dimensional confidence ellipsoid for $\bm{\beta}$ within the corresponding subspace. The union of all such $d_{K}$-dimensional confidence ellipsoids creates one confidence set for $\bm{\beta}$, coming from offline uncertainty. Let’s call it $\mathcal{C}^{t}(\beta,\operatorname{off})$. On the other hand, we also have the purely online $d_{A}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\beta,\operatorname{on})$. We can thus sharpen our optimism by being optimistic over the smaller intersection $\mathcal{C}^{t}(\beta,\operatorname{off})\cap\mathcal{C}^{t}(\beta,\operatorname{on})$ instead of just over $\mathcal{C}^{t}(\beta,\operatorname{on})$ – thereby performing "more localized" optimism than LinUCB. This is the idea behind Algorithm 2, which we thus call LOCAL-UCB. Figure 1 illustrates different ways in which taking this intersection can affect optimism.
We formalize this intuition in Algorithm 2 by formulating the sharpened optimism as an optimization problem in Line 4. The first constraint respresents the low dimensional confidence ellipsoids in the subspace spanned by a given U, while the third constraint lets U range over all of $\mathcal{C}_{U,\operatorname{off}}$. The second constraint merely represents the usual high dimensional ellipsoid.

### v2

@@ 3 Estimating Latent Subspaces Offline
To minimize the regret, one must learn the reward parameter $\bm{\beta}_{\star}$ online. However, it is much easier to search among all latent states $\bm{\theta}\in\mathbb{R}^{d_{K}}$ than to search among all possible reward parameters $\bm{\beta}\in\mathbb{R}^{d_{A}}$ since typically, $d_{K}\ll d_{A}$. So, it will help to learn some projection matrix $\hat{\textbf{U}}^{\top}\approx\textbf{U}_{*}^{\top}:\mathbb{R}^{d_{A}}\to\mathbb{R}^{d_{K}}$ offline so that for any estimate $\hat{\bm{\beta}}_{t}$ of $\bm{\beta}_{\star}$, $\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{t}$ is an estimate of $\hat{\bm{\theta}}_{t}\in\mathbb{R}^{d_{K}}$. This amounts to learning a subspace of the feature space from logged bandit data. We therefore provide a method for Subspace estimation from Offline Latent bandit Data (SOLD) in Algorithm 1. Recall that since the learnt subspace is approximate, we also need to compute the uncertainty over the subspace to get a subspace confidence set that we can use online.
@@
3:Estimate reward parameters $\hat{\bm{\beta}}_{n,i}\leftarrow\textbf{V}_{n,i}^{-1}\textbf{b}_{n,i}$, where $\textbf{V}_{n,i}\leftarrow\mu\textbf{I}+\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)\phi(x,a)^{\top}$ and $\textbf{b}_{n,i}\leftarrow\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)r$ for $i=1,2$.
4:Compute $\textbf{M}_{n}\leftarrow\frac{1}{2}(\hat{\bm{\beta}}_{n,1}\hat{\bm{\beta}}_{n,2}^{\top}+\hat{\bm{\beta}}_{n,2}\hat{\bm{\beta}}_{n,1}^{\top})$ and compute $\overline{\textbf{M}}_{N}\leftarrow\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}$.
5:Compute $\overline{\textbf{D}}_{N,i}\leftarrow\frac{1}{N}\sum_{n=1}^{N}(I-\mu\textbf{V}_{n,i}^{-1})$, $i=1,2$.
6:Obtain $\hat{\textbf{U}}$, the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$.
@@
To extract the $d_{K}$-dimensional subspace, we aim to estimate $\mathbb{E}[\bm{\beta}\bm{\beta}^{\top}]=\textbf{U}_{\star}\mathbb{E}[\bm{\theta}\bm{\theta}^{\top}]\textbf{U}_{\star}^{\top}$, which has the same $d_{K}$-dimensional span as $\textbf{U}_{\star}$. This cannot be achieved by using a single estimator $\hat{\bm{\beta}}_{n}$ for each trajectory $\tau_{n}$ and averaging the outer products $\hat{\bm{\beta}}_{n}\hat{\bm{\beta}}_{n}^{\top}$ across all $n$. That is because the per-reward noise $\epsilon$ will be shared by both copies of $\hat{\bm{\beta}}_{n}$, and so the variance of $\epsilon$ will make $\mathbb{E}[\hat{\bm{\beta}}_{n}\hat{\bm{\beta}}_{n}^{\top}]$ full rank. We therefore split each trajectory $\tau_{n}$ to obtain two independent estimators $\hat{\bm{\beta}}_{n,1},\hat{\bm{\beta}}_{n,2}$, compute the outer products $\hat{\bm{\beta}}_{n,1}^{\top}\hat{\bm{\beta}}_{n,2}$, and obtain the top $d_{K}$ eigenvectors of the mean outer product across trajectories.
However, there is a further wrinkle here. We cannot simply take the top $d_{K}$ eigenvectors of the mean outer product $\overline{\textbf{M}}_{N}$. One can compute that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}\textbf{U}_{\star}\bm{\theta}_{n}\bm{\theta}_{n}^{\top}\textbf{U}_{\star}^{\top}\textbf{D}_{n,2}]$. To separate $\mathbb{E}[\bm{\beta}_{n}\bm{\beta}_{n}^{\top}]$ from this, we need $\textbf{D}_{n,1},\bm{\beta}_{n},\textbf{D}_{n,2}$ to be independent. If $\pi_{b}$ does not use $\bm{\theta}$ and contexts are generated independently of each other and of $\bm{\theta}$, then this is satisfied. Intuitively, we need the offline trajectories to be non-adaptive. In fact, we show in the lemma below that if any of these three conditions is violated, then it is in fact impossible to determine the latent subspace $\textbf{U}^{\star}$ using any method, even with infinitely many infinitely long trajectories.
@@
there exist two different linear latent contextual bandits with orthogonal latent subspaces satisfying the condition, and a behavior policy $\pi_{b}$ so that the offline data distributions are indistinguishable and cover all $(x,a)$ pairs with probability at least $1/4$. Since the latent subspaces are orthogonal, an action that gives the maximum reward on one latent bandit gives reward 0 on the other.
@@
This is satisfied when the offline data comes from randomized controlled trials or A/B testing, which are common sources of offline datasets. Even if this is not satisfied, Algorithm 1 can learn a good subspace whenever $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ has eigenspace close to the span of $\textbf{U}_{\star}$, e.g. when high-reward actions contribute heavily to $\textbf{D}_{n,i}$. This can happen if offline trajectories were collected to maximize rewards.
Returning to our scrutiny of $\overline{\textbf{M}}_{N}$, let the covariance matrix of $\bm{\theta}$ be $\Lambda$ and let its mean be $\mu_{\theta}$. Then we have that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}]\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}\mathbb{E}[\textbf{D}_{n,2}]$. So, we still cannot merely consider the top $d_{K}$ eigenvectors of $\overline{\textbf{M}}_{N}$ without accounting for $\textbf{D}_{n,1}$. Intuitively, $\textbf{D}_{n,1}$ captures the distortion in reward estimation caused by regularization in ridge regression [^2] This is not unique to regularization. Pseudo-inverses cause an analogous problem of distortion caused by unseen actions.. We therefore construct correction matrices $\overline{\textbf{D}}_{N,i}$ and use them to "neutralize" the distortion from regularization. In particular, $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is an estimator for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Crucially, this allows us to aggregate information across many trajectories to overcome the challenge of learning from short trajectories. We can now take the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to estimate the subspace determined by $\textbf{U}_{\star}$. To give guarantees, we must make a coverage assumption. Unlike in standard offline RL, where only coverage along actions is needed, we also need coverage along latent states.
@@
Rewards $|r_{n,h}|\leq R$  [^3] RR-bounded rewards are automatically $R$-subgaussian. We can easily extend our results to more general subgaussian rewards, but stick to bounded rewards for simplicity of proofs. for all $n,h$, $\|\phi(x,a)\|_{2}\leq 1$ and $\|\bm{\beta}\|_{2}\leq R$. Also, $\lambda_{A}:=\min_{i=1,2}\lambda_{\min}(\mathbb{E}[\textbf{D}_{n,i}])>0$ and $\lambda_{\theta}:=\frac{1}{R^{2}}\lambda_{\min}(\Lambda)>0$.
Intuitively, $\lambda_{A}$ measures coverage along actions, while $\lambda_{\theta}$ measures coverage along latent states $\bm{\theta}$. Both must be non-zero to expect satisfactory estimation of the subspace. We can then use confidence bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ to give a data-dependent confidence bounds $\Delta_{\operatorname{off}}$ for the projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, as in Theorem 1 below. In one instantiation, Propositions 1 and 2 in Appendix C derive simple data-dependent bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ respectively. Under this choice, we control the growth of $\Delta_{\operatorname{off}}$ in terms of the unknown problem parameters at the end of Theorem 1.
@@
Let $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ and $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}\leq\Delta_{D}$ for $i=1,2$ with probability $1-\delta/3$ each. Then, with probability $1-\delta$, $\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}\leq\Delta_{\operatorname{off}}$, where for $B_{D}=\|\overline{\textbf{D}}_{N}^{-1}\|_{2}$ and $\hat{\lambda}:=\lambda_{d_{K}}(\overline{\textbf{M}}_{N})-\lambda_{d_{K}+1}(\overline{\textbf{M}}_{N})$,
$\Delta_{\operatorname{off}}$ $=\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}(R^{2}+\Delta_{M})\Delta_{D}\right$.
$\left.\qquad+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}\right).$
@@
Estimating $d_{K}$ offline.
As our estimator $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is approximately rank-$d_{K}$, the number of nonzero eigenvalues of the estimator is a principled heuristic for determining $d_{K}$.
@@
Naively performing PCA on the raw rewards or on single reward estimates $\hat{\beta}_{n}$ can lead to erroneous subspaces – as while the PCA target is linear-algebraically similar to $\overline{\textbf{M}}_{N}$, it is statistically different. The PCA target (e.g. $\mathbb{E}[\hat{\beta}_{n}\hat{\beta}_{n}^{\top}]$) is typically full rank due to the variance of the per-reward noise $\epsilon$. On the other hand, PMF Mnih and Salakhutdinov, 2007a offers neither confidence bounds on the estimated subspace, nor a principled method for determining $d_{K}$.
4 Offline Data Sharpens Online Optimism
Here, we motivate and describe LOCAL-UCB, a natural algorithm that accelerates LinUCB with offline data. The core idea is sharpening optimism by being optimistic over the intersection of two confidence sets – one obtained using offline and online data and another purely from online data.
We geometrically motivate our update rule here, and illustrate it in Figure 1. After any $t$ steps, we can construct a $d_{K}$-dimensional confidence ellipsoid for every subspace in the subspace confidence set obtained from SOLD. The union of all these ellipsoids gives us our "offline confidence set" [^4] The set is not only dependent on offline data, since online data is used to construct the $d_{K}$-dimensional ellipsoids., called $\mathcal{C}^{t}_{\operatorname{off}}(\bm{\beta})$. The usual $d_{A}$-dimensional ellipsoid forms our "online confidence set." We call this $\mathcal{C}^{t}_{\operatorname{on}}(\bm{\beta})$. Since the true parameter lies in both confidence sets with high probability, being optimistic over their intersection allows us to sharpen or "further localize" optimism. Even though the offline confidence set uses both offline and online data, it will never shrink to a point due to the frozen subspace confidence set. So, we need the intersection of both sets to be sharply optimistic. This is the intuition behind LOCAL-UCB.
We formalize this intuition in Algorithm 2 by formulating the sharpened optimism as an optimization problem in step 4. The first two constraints represent the low dimensional confidence ellipsoids in the subspace spanned by a given U, while the next two merely represent the usual high dimensional ellipsoid. The remaining constraints let U range over our subspace confidence set.